# Unfitted hybrid high-order methods stabilized by polynomial extension for elliptic interface problems

### v1

@@ Introduction
where we recall that the superscript + is used to indicate the extension of a polynomial originally defined on $T^{i}$ to $T^{i}\cup\bigcup_{S\in\mathcal{N}_{i}^{-1}(T)}S^{i}$.
@@
$\boldsymbol{\delta}_{T,S,i}(v)$ :=$:=\big{\{}\boldsymbol{G}_{T^{i}}^{k}\big{(}\hat{I}_{T}^{k}(v)^{\mathcal{N}}\big{)}+\delta_{i1}\boldsymbol{L}_{T^{1}}^{k}(\llbracket v\rrbracket_{\Gamma})\big{\}}^{+}|_{S^{i}}-\nabla v_{i}|_{S^{i}},\qquad\forall S\in\mathcal{N}_{i}^{-1}(T),$ (50b)
and recall the error measure $\epsilon_{T,i}(v)$ defined in (40). Then, the following holds:
$\|\boldsymbol{\delta}_{T,i}(v)\|_{T^{i}}+h_{T}^{\frac{1}{2}}\|\boldsymbol{\delta}_{T,i}(v)\|_{(\partial T)^{i}\cup T^{\Gamma}}$ $\lesssim h_{T}^{-1}\big{\{}\epsilon_{T,i}(v_{i})+\delta_{i1}\epsilon_{T,2}(v_{2})\big{\}},$ (51a)
$\sum_{S\in\mathcal{N}_{i}^{-1}(T)}\Big{\{}\|\boldsymbol{\delta}_{T,S,i}(v)\|_{S^{i}}+h_{S}^{\frac{1}{2}}\|\boldsymbol{\delta}_{T,S,i}(v)\|_{(\partial S)^{i}\cup S^{\Gamma}}\Big{\}}$ $\lesssim h_{T}^{-1}\big{\{}\epsilon_{T,i}(v_{i})+\delta_{i1}\epsilon_{T,2}(v_{2})\big{\}}.$ (51b)
@@
[…]
@@
Recalling the definition (40) of $\epsilon_{T,i}(v)$ gives
$\|\boldsymbol{\delta}^{\prime}_{T,i}(v)\|_{T^{i}}\lesssim h_{T}^{-1}\big{\{}\epsilon_{T,i}(v_{i})+\delta_{i1}\epsilon_{T,2}(v_{2})\big{\}},$ (52)
@@
$\|\boldsymbol{\delta}^{\prime}_{T,i}(v)\|_{T^{i}}+h_{T}^{\frac{1}{2}}\|\boldsymbol{\delta}^{\prime}_{T,i}(v)\|_{(\partial T)^{i}\cup T^{\Gamma}}\lesssim\|\boldsymbol{\delta}^{\prime}_{T,i}(v)\|_{T^{i}}\lesssim h_{T}^{-1}\big{\{}\epsilon_{T,i}(v_{i})+\delta_{i1}\epsilon_{T,2}(v_{2})\big{\}}.$
Finally, since $\boldsymbol{\delta}_{T,i}(v)=\boldsymbol{\delta}^{\prime}_{T,i}(v)-\nabla(v_{i}-I_{T^{i}}^{k+1}(v_{i}))$ invoking the triangle inequality proves (51a).
@@
$\|\boldsymbol{\delta}^{\prime}_{T,i}(v)^{+}\|_{S^{i}}+h_{S}^{\frac{1}{2}}\|\boldsymbol{\delta}^{\prime}_{T,i}(v)^{+}\|_{(\partial S)^{i}\cup S^{\Gamma}}\lesssim\|\boldsymbol{\delta}^{\prime}_{T,i}(v)\|_{T^{i}}\lesssim h_{T}^{-1}\big{\{}\epsilon_{T,i}(v_{i})+\delta_{i1}\epsilon_{T,2}(v_{2})\big{\}}.$
Invoking the triangle inequality and recalling (40) proves (51b). ∎
@@
Let $u$ be the weak solution to (2). Assume that $u\in H^{s}(\Omega_{1}\cup\Omega_{2})$, $s>\frac{3}{2}$. Set $\Phi_{h}(\hat{w}_{h}):=a_{h}(\hat{I}_{h}^{k}(u),\hat{w}_{h})-\ell_{h}(\hat{w}_{h})$ for all $\hat{w}_{h}\in\widehat{\mathcal{U}}_{h0}$. The following holds:
$|\Phi_{h}(\hat{w}_{h})|\lesssim{}$$\Bigg{\{}\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}h_{T}^{-2}\epsilon_{T,i}(u_{i})^{2}\Bigg{\}}^{\frac{1}{2}}\|\hat{w}_{h}\|_{h0}.$ (53)
@@
$A_{1}={}$ $\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}(\boldsymbol{\delta}_{T,i}(u),\nabla w_{T^{i}})_{T^{i}},$
$A_{2}={}$ $\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}\bigg{\{}(\boldsymbol{\delta}_{T,i}(u){\cdot}\boldsymbol{n}_{T},w_{(\partial T)^{i}}-w_{T^{i}})_{(\partial T)^{i}}-\delta_{i1}(\boldsymbol{\delta}_{T,1}(u){\cdot}\boldsymbol{n}_{\Gamma},\llbracket w_{T}\rrbracket_{\Gamma})_{T^{\Gamma}}$
$+\sum_{S\in\mathcal{N}_{i}^{-1}(T)}\big{\{}(\boldsymbol{\delta}_{T,S,i}(u){\cdot}\boldsymbol{n}_{S},w_{(\partial S)^{i}}-w_{S^{i}})_{(\partial S)^{i}}-\delta_{i1}(\boldsymbol{\delta}_{T,S,1}(u){\cdot}\boldsymbol{n}_{\Gamma},\llbracket w_{S}\rrbracket_{\Gamma})_{S^{\Gamma}}\big{\}}\bigg{\}},$
$A_{3}={}$ $s_{h}^{\circ}(\hat{I}_{h}^{k}(u),\hat{w}_{h})+\sum_{T\in\mathcal{T}_{h}^{\textup{cut}}}\kappa_{1}h_{T}^{-1}(\llbracket I_{T}^{k+1}(u)-u\rrbracket_{\Gamma},\llbracket w_{T}\rrbracket_{\Gamma})_{T^{\Gamma}}+s_{h}^{\mathcal{N}}(\hat{I}_{h}^{k}(u),\hat{w}_{h}),$
where $\boldsymbol{\delta}_{T,i}(u)$ and $\boldsymbol{\delta}_{T,S,i}(u)$ are defined in (50) for all $(T,i)\in\mathcal{P}_{h}^{\textup{OK}}$, whereas we set $\boldsymbol{\delta}_{T,i}(u):=\nabla(I^{k+1}_{T^{i}}(u_{i})-u_{i})$ for all $(T,i)\in\mathcal{P}_{h}^{\textup{KO}}$, and where $\llbracket I_{T}^{k+1}(u)-u\rrbracket_{\Gamma}$ is defined in Lemma 4.2. Owing to (49), we infer that the term $A_{1}$ can be rewritten as follows:
$A_{1}=\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}\bigg{\{}(\boldsymbol{\delta}_{T,i}(u),\nabla w_{T^{i}})_{T^{i}}+\sum_{S\in\mathcal{N}_{i}^{-1}(T)}(\boldsymbol{\delta}_{T,S,i}(u),\nabla w_{S^{i}})_{S^{i}}\bigg{\}}.$
@@
$|A_{1}|\lesssim\Bigg{\{}\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}h_{T}^{-2}\big{\{}\epsilon_{T,i}(u_{i})+\delta_{i1}\epsilon_{T,2}(u_{2})\big{\}}^{2}\Bigg{\}}^{\frac{1}{2}}\Bigg{\{}\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}\|\nabla w_{T^{i}}\|_{T^{i}}^{2}\Bigg{\}}^{\frac{1}{2}}.$
Since $\kappa_{1}\leq\kappa_{2}$ and recalling the definition (42) of the stability norm $\|{\cdot}\|_{h0}$, we infer that
$|A_{1}|\lesssim\Bigg{\{}\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}h_{T}^{-2}\epsilon_{T,i}(u_{i})^{2}\Bigg{\}}^{\frac{1}{2}}\|\hat{w}_{h}\|_{h0}.$
The same arguments prove that
$|A_{2}|\lesssim\Bigg{\{}\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}h_{T}^{-2}\epsilon_{T,i}(u_{i})^{2}\Bigg{\}}^{\frac{1}{2}}\|\hat{w}_{h}\|_{h0}.$
@@
$|A_{31}|$$\leq\Bigg{\{}\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}h_{T}^{-1}\|\Pi_{(\partial T)^{i}}^{k}(u_{i}-I_{T^{i}}^{k+1}(u_{i}))\|_{(\partial T)^{i}}^{2}\Bigg{\}}^{\frac{1}{2}}|\hat{w}_{h}|_{\textsc{s}}$
$\leq\Bigg{\{}\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}h_{T}^{-1}\|u_{i}-I_{T^{i}}^{k+1}(u_{i})\|_{(\partial T)^{i}}^{2}\Bigg{\}}^{\frac{1}{2}}|\hat{w}_{h}|_{\textsc{s}},$
@@
$|A_{31}|\leq\Bigg{\{}\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\sum_{S\in\{T\}\cup\mathcal{N}_{i}^{-1}(T)}\kappa_{i}h_{S}^{-1}\|u_{i}-I_{T^{i}}^{k+1}(u_{i})^{+}\|_{(\partial S)^{i}}^{2}\Bigg{\}}^{\frac{1}{2}}|\hat{w}_{h}|_{\textsc{s}}.$
@@
$|A_{31}|\leq\Bigg{\{}\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}h_{T}^{-2}\epsilon_{T,i}(u_{i})^{2}\Bigg{\}}^{\frac{1}{2}}|\hat{w}_{h}|_{\textsc{s}}.$
@@
$|A_{32}|\leq\Bigg{\{}\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}h_{T}^{-2}\epsilon_{T,i}(u_{i})^{2}\Bigg{\}}^{\frac{1}{2}}s_{h}^{\Gamma}(\hat{w}_{h},\hat{w}_{h})^{\frac{1}{2}}.$
Finally, $s_{h}^{\mathcal{N}}(\hat{I}_{h}^{k}(u),\hat{w}_{h})=0$ owing to (49). Putting everything together proves the claim. ∎
@@
Let $u$ be the weak solution to (2). Assume that $u\in H^{s}(\Omega_{1}\cup\Omega_{2})$ with $s\in(\frac{3}{2},k+2]$. The following holds:
$\Bigg{\{}\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}\|\nabla(u_{i}-u_{T^{i}})\|_{T^{i}}^{2}\Bigg{\}}^{\frac{1}{2}}\lesssim\Bigg{\{}\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}h_{T}^{2(s-1)}|E_{i}^{s}(u_{i})|_{H^{s}(\Delta_{2}(T))}^{2}\Bigg{\}}^{\frac{1}{2}}\lesssim h^{s-1}\sum_{i\in\{1,2\}}\kappa_{i}^{\frac{1}{2}}|u_{i}|_{H^{s}(\Omega_{i})}.$ (54)

### v2

@@ Introduction
where we recall that the superscript + is used to indicate the extension of a polynomial originally defined on $T^{i}$ to $\Delta(T)$.
@@
$\boldsymbol{\delta}_{T,S,i}(v)$ $:=\big\{\boldsymbol{G}_{T^{i}}^{k}\big(\hat{I}_{T}^{k}(v)^{\mathcal{N}}\big)+\delta_{i1}\boldsymbol{L}_{T^{1}}^{k}(\llbracket v\rrbracket_{\Gamma})\big\}^{+}|_{S^{i}}-\nabla v_{i}|_{S^{i}},\qquad\forall S\in\mathcal{N}_{i}^{-1}(T),$ (50b)
and recall the error measures $\epsilon_{T,i}(v_{i})$ and $\epsilon^{\Gamma}_{T,i}(v)$ defined in (40a)-(40b). Then, the following holds:
$\|\boldsymbol{\delta}_{T,i}(v)\|_{T^{i}}+h_{T}^{\frac{1}{2}}\|\boldsymbol{\delta}_{T,i}(v)\|_{(\partial T)^{i}\cup T^{\Gamma}}$ $\lesssim h_{T}^{-1}\big\{\epsilon_{T,i}(v_{i})+\delta_{i1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon^{\Gamma}_{T,i}(v)}\big\},$ (51a)
$\sum_{S\in\mathcal{N}_{i}^{-1}(T)}\Big\{\|\boldsymbol{\delta}_{T,S,i}(v)\|_{S^{i}}+h_{S}^{\frac{1}{2}}\|\boldsymbol{\delta}_{T,S,i}(v)\|_{(\partial S)^{i}\cup S^{\Gamma}}\Big\}$ $\lesssim h_{T}^{-1}\big\{\epsilon_{T,i}(v_{i})+\delta_{i1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon^{\Gamma}_{T,i}(v)}\big\}.$ (51b)
@@
$={}$ $(\boldsymbol{G}_{T^{i}}^{k}\big(\hat{I}_{T}^{k}(v)^{\mathcal{N}}\big)+\delta_{i1}\boldsymbol{L}_{T^{1}}^{k}(\llbracket I_{\mathcal{T}_{h}}^{k+1}(v)\rrbracket_{\Gamma})-\nabla I_{T^{i}}^{k+1}(v_{i}),\boldsymbol{\delta}^{\prime}_{T,i}(v))_{T^{i}}$
@@
Recalling the definition (40a) of $\epsilon_{T,i}(v)$ and the definition (40b) of $\epsilon^{\Gamma}_{T,i}(v)$ gives
$\|\boldsymbol{\delta}^{\prime}_{T,i}(v)\|_{T^{i}}\lesssim h_{T}^{-1}\big\{\epsilon_{T,i}(v_{i})+\delta_{i1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon^{\Gamma}_{T,i}(v)}\big\},$ (52)
@@
$\|\boldsymbol{\delta}^{\prime}_{T,i}(v)\|_{T^{i}}+h_{T}^{\frac{1}{2}}\|\boldsymbol{\delta}^{\prime}_{T,i}(v)\|_{(\partial T)^{i}\cup T^{\Gamma}}\lesssim\|\boldsymbol{\delta}^{\prime}_{T,i}(v)\|_{T^{i}}\lesssim h_{T}^{-1}\big\{\epsilon_{T,i}(v_{i})+\delta_{i1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon^{\Gamma}_{T,i}(v)}\big\}.$
Finally, since $\boldsymbol{\delta}_{T,i}(v)=\boldsymbol{\delta}^{\prime}_{T,i}(v)-\nabla(v_{i}-I_{T^{i}}^{k+1}(v_{i}))$, invoking the triangle inequality proves (51a).
@@
$\|\boldsymbol{\delta}^{\prime}_{T,i}(v)^{+}\|_{S^{i}}+h_{S}^{\frac{1}{2}}\|\boldsymbol{\delta}^{\prime}_{T,i}(v)^{+}\|_{(\partial S)^{i}\cup S^{\Gamma}}\lesssim\|\boldsymbol{\delta}^{\prime}_{T,i}(v)\|_{T^{i}}\lesssim h_{T}^{-1}\big\{\epsilon_{T,i}(v_{i})+\delta_{i1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon^{\Gamma}_{T,i}(v)}\big\}.$
Invoking the triangle inequality and recalling (40a) proves (51b). ∎
@@
Let $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}$} be the weak solution to (2). Assume that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}\in H^{s}(\Omega_{1}\cup\Omega_{2})$, $s>\frac{3}{2}$. Set $\Phi_{h}(\hat{w}_{h}):=a_{h}(\hat{I}_{h}^{k}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}),\hat{w}_{h})-\ell_{h}(\hat{w}_{h})$ for all $\hat{w}_{h}\in\widehat{\mathcal{U}}_{h0}$. The following holds:
$|\Phi_{h}(\hat{w}_{h})|\lesssim{}$ $\Bigg\{\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}h_{T}^{-2}\big\{\epsilon_{T,i}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}})+\delta_{i1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon^{\Gamma}_{T,i}(u^{\textup{ex}})}\big\}^{2}\Bigg\}^{\frac{1}{2}}\|\hat{w}_{h}\|_{h0}.$ (53)
@@
$A_{1}={}$ $\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}(\boldsymbol{\delta}_{T,i}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}),\nabla w_{T^{i}})_{T^{i}},$
$A_{2}={}$ $\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}\bigg\{(\boldsymbol{\delta}_{T,i}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}){\cdot}\boldsymbol{n}_{T},w_{(\partial T)^{i}}-w_{T^{i}})_{(\partial T)^{i}}-\delta_{i1}(\boldsymbol{\delta}_{T,1}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}){\cdot}\boldsymbol{n}_{\Gamma},\llbracket w_{T}\rrbracket_{\Gamma})_{T^{\Gamma}}$
$+\sum_{S\in\mathcal{N}_{i}^{-1}(T)}\big\{(\boldsymbol{\delta}_{T,S,i}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}){\cdot}\boldsymbol{n}_{S},w_{(\partial S)^{i}}-w_{S^{i}})_{(\partial S)^{i}}-\delta_{i1}(\boldsymbol{\delta}_{T,S,1}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}){\cdot}\boldsymbol{n}_{\Gamma},\llbracket w_{S}\rrbracket_{\Gamma})_{S^{\Gamma}}\big\}\bigg\},$
$A_{3}={}$ $s_{h}^{\circ}(\hat{I}_{h}^{k}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}),\hat{w}_{h})+\sum_{T\in\mathcal{T}_{h}^{\textup{cut}}}\kappa_{1}h_{T}^{-1}(\llbracket I_{T}^{k+1}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}\rrbracket_{\Gamma},\llbracket w_{T}\rrbracket_{\Gamma})_{T^{\Gamma}}+s_{h}^{\mathcal{N}}(\hat{I}_{h}^{k}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}),\hat{w}_{h}),$
where $\boldsymbol{\delta}_{T,i}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}})$ and $\boldsymbol{\delta}_{T,S,i}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}})$ are defined in (50) for all $(T,i)\in\mathcal{P}_{h}^{\textup{OK}}$, whereas we set $\boldsymbol{\delta}_{T,i}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}):=\nabla(I^{k+1}_{T^{i}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}})$ for all $(T,i)\in\mathcal{P}_{h}^{\textup{KO}}$, and where $\llbracket I_{T}^{k+1}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}\rrbracket_{\Gamma}$ is defined in Lemma 4.3. Owing to (49), we infer that the term $A_{1}$ can be rewritten as follows:
$A_{1}=\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}\bigg\{(\boldsymbol{\delta}_{T,i}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}),\nabla w_{T^{i}})_{T^{i}}+\sum_{S\in\mathcal{N}_{i}^{-1}(T)}(\boldsymbol{\delta}_{T,S,i}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}),\nabla w_{S^{i}})_{S^{i}}\bigg\}.$
@@
$|A_{1}|$ $\lesssim\Bigg\{\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}h_{T}^{-2}\big\{\epsilon_{T,i}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}})+\delta_{i1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon^{\Gamma}_{T,i}(u^{\textup{ex}})}\big\}^{2}\Bigg\}^{\frac{1}{2}}\Bigg\{\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}\|\nabla w_{T^{i}}\|_{T^{i}}^{2}\Bigg\}^{\frac{1}{2}}$
$\leq\Bigg\{\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}h_{T}^{-2}\big\{\epsilon_{T,i}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}})+\delta_{i1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon^{\Gamma}_{T,i}(u^{\textup{ex}})}\big\}^{2}\Bigg\}^{\frac{1}{2}}\|\hat{w}_{h}\|_{h0},$
where the last bound follows from the definition (42) of the stability norm $\|{\cdot}\|_{h0}$. The same arguments prove that
$|A_{2}|\lesssim\Bigg\{\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}h_{T}^{-2}\big\{\epsilon_{T,i}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}})+\delta_{i1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon^{\Gamma}_{T,i}(u^{\textup{ex}})}\big\}^{2}\Bigg\}^{\frac{1}{2}}\|\hat{w}_{h}\|_{h0}.$
@@
$|A_{31}|$ $\leq\Bigg\{\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}h_{T}^{-1}\|\Pi_{(\partial T)^{i}}^{k}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}-I_{T^{i}}^{k+1}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}))\|_{(\partial T)^{i}}^{2}\Bigg\}^{\frac{1}{2}}|\hat{w}_{h}|_{\textsc{s}}$
$\leq\Bigg\{\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}h_{T}^{-1}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}-I_{T^{i}}^{k+1}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}})\|_{(\partial T)^{i}}^{2}\Bigg\}^{\frac{1}{2}}|\hat{w}_{h}|_{\textsc{s}},$
@@
$|A_{31}|\leq\Bigg\{\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\sum_{S\in\{T\}\cup\mathcal{N}_{i}^{-1}(T)}\kappa_{i}h_{S}^{-1}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}-I_{T^{i}}^{k+1}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}})^{+}\|_{(\partial S)^{i}}^{2}\Bigg\}^{\frac{1}{2}}|\hat{w}_{h}|_{\textsc{s}}.$
@@
$|A_{31}|\leq\Bigg\{\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}h_{T}^{-2}\epsilon_{T,i}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}})^{2}\Bigg\}^{\frac{1}{2}}|\hat{w}_{h}|_{\textsc{s}}.$
@@
$|A_{32}|\leq\Bigg\{\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}h_{T}^{-2}\big\{\epsilon_{T,i}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}})+\delta_{i1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon^{\Gamma}_{T,i}(u^{\textup{ex}})}\big\}^{2}\Bigg\}^{\frac{1}{2}}s_{h}^{\Gamma}(\hat{w}_{h},\hat{w}_{h})^{\frac{1}{2}}.$
Finally, $s_{h}^{\mathcal{N}}(\hat{I}_{h}^{k}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}),\hat{w}_{h})=0$ owing to (49). Putting everything together proves the claim. ∎
@@
Let $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}$} be the weak solution to (2). Assume that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}\in H^{s}(\Omega_{1}\cup\Omega_{2})$ with $s\in(\frac{3}{2},k+2]$. The following holds:
$\Bigg\{\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}\|\nabla({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}-u_{T^{i}})\|_{T^{i}}^{2}\Bigg\}^{\frac{1}{2}}\lesssim\Bigg\{\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}h_{T}^{2(s-1)}|E_{i}^{s}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}})|_{H^{s}(\Delta_{2}(T))}^{2}\Bigg\}^{\frac{1}{2}}\lesssim h^{s-1}\sum_{i\in\{1,2\}}\kappa_{i}^{\frac{1}{2}}|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}|_{H^{s}(\Omega_{i})}.$ (54)